# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v2

@@ Appendix H Evaluating Dataset Auditing in the Fine-tuning Scenario
Results. The outcomes of applying evasion attacks in the fine-tuning scenario are presented in Table 8. The results indicate that dataset auditing methods exhibit vulnerabilities similar to those observed when training from scratch. Both IF-based methods and EF-based methods frequently fail when confronted with evasion attacks such as DP-SGD, RS, or Reprogramming. This suggests that the inherent weaknesses of current auditing techniques are not limited to models trained from scratch but also extend to the fine-tuning paradigm.
Appendix I Forgery Attacks in the Black-box Scenario
This section evaluates forgery attacks when the adversary only possesses black-box access to the target model. Lacking parameter access hinders gradient-based attacks. Thus, the adversary may distill a substitute model from the target model using an auxiliary dataset. Then they can conduct forgery attacks on this substitute model, relying on the transferability of adversarial attacks to affect the original target model.
Experimental Settings. We simulate the black-box setting where the target model is the ResNet-18 and the substitute model employs the MobileNetV2 architecture[72]. All other settings remain consistent with the main experiments.
Results. From Table 9, the success rate of forgery attacks generally decreases in the black-box setting compared to the white-box scenario. However, despite the reduction, most dataset auditing methods ($e.g.$, DUA, DVBW, and UBW) remain susceptible to forgery even under these stricter black-box constraints. This indicates that even without direct model access, achieving dataset forgery remains a practical threat against a significant portion of current auditing techniques.
Appendix J Comparison with Other Related Data Protection Techniques
Data protection has long been a critical aspect of information management [63]. In the last several decades, researchers have developed many different techniques to achieve various requirements of data protection, as follows.
Encryption. Encryption-based techniques aim to transform the sensitive data into unreadable ciphertext, and any unauthorized party without the secret key cannot access the data [5, 27]. Advancements in encryption-based techniques like homomorphic encryption (HE) [65, 15] now allow computations on encrypted data, making it promising to be adopted in DL systems. However, encryption-based methods still suffer from the limitations of excessive latency and resource demands.
Differential Privacy. Differential privacy (DP) ensures that the inclusion or exclusion of any single data point in a dataset does not significantly affect the output [24]. This is typically achieved by injecting controlled noise into data, queries, or model training processes [24, 1]. While DP provides strong, mathematically provable privacy guarantees, its application in DL introduces significant challenges. For instance, the added noise often degrades model utility. The trade-off between privacy and utility makes DP a powerful but costly solution for privacy-preserving DL [1, 85].
Traditional Data Watermarking. Traditional data watermarking techniques help track ownership and usage of data by embedding hidden markers ($i.e.$, watermark) into the data [44, 87]. The data owner can extract the watermark from the data and prove their ownership. Multiple solutions to embedding a watermark into data have been proposed, such as embedding watermarks into the frequency domain or utilizing DL models [9, 4]. Recently, some works have also explored using data watermarking to mark AI-generated content and models [86, 75, 45, 70]. However, traditional data watermarking techniques may not be feasible for dataset auditing since the watermark in the data can hardly be inherited by the DL models, and their owners will not disclose their training details (especially their training data).
Unlearnable Data. Unlearnable data techniques aim to add perturbations to the data so that the DL model cannot learn from it during training [40, 81]. The model trained on the unlearnable dataset will have a low utility. However, in the scenario of protecting a public dataset where we intend to prevent unauthorized usage instead of making it completely useless, these techniques may not be feasible.
In general, previous data protection techniques either have a significant impact on the data and model utility ($e.g.$, HE, DP, and unlearnable data) or are not feasible ($e.g.$, traditional data watermarking) in DL scenarios [51, 33]. Consequently, dataset auditing, as a technique that has negligible influence on the data utility and can protect the privacy and copyright of data owners, has become an emerging yet important and promising solution to data protection [48, 20].

### v3

@@ 5.4 Evaluating Dataset Auditing in the Fine-tuning Scenario
The experimental results of applying evasion attacks in the fine-tuning scenario are presented in Table 6. The results indicate that dataset auditing methods exhibit vulnerabilities similar to those observed when training from scratch. Both IF-based methods and EF-based methods frequently fail when confronted with evasion attacks such as DP-SGD, RS, or Reprogramming. This suggests that the inherent weaknesses of existing dataset auditing techniques are not limited to models trained from scratch but also extend to the fine-tuning paradigm.
5.5 Results of Forgery Attacks
For forgery attacks, we assume that the adversary wants to make their own dataset ($i.e.$, the attack auxiliary dataset) to be falsely identified as the ‘trained dataset’ of other independent models. We use the model trained on the unmarked release dataset as the independent model. The results of white-box forgery attacks are in Table 7, and the experiments in the black-box scenario can be found in Table 8.
Figure 3: Relative time overhead of different evasion attacks. The overhead is compared with the baseline time without any specific evasion attack ($i.e.$, ‘No Attack’ with a relative overhead normalized to 1.0).
Takeaway 5: Existing dataset auditing methods are overall vulnerable to forgery attacks.
Takeaway 6: Dataset auditing methods that involve more complex decision-making and auditing processes are more resistant to forgery.
5.6 Attack Overhead
5.7 Summary of Evaluated Dataset Auditing Methods
Our analysis reveals that no single method excels across all 5 dimensions. The most critical finding is the pervasive lack of adversarial robustness and distinctiveness, as also analyzed in the previous section. On the other hand, fidelity is typically high, indicating minimal impact on model utility in benign settings. A clear divide exists in stealthiness. IF-based methods inherently excel in stealthiness, while EF-based methods depend on the embedded feature. Efficiency spans a wide spectrum, from computationally cheap methods to highly demanding ones requiring intensive processing overhead.
6 Discussion
6.1 Related Works
Since dataset auditing in DL is an emerging topic, there are only a limited number of surveys or benchmarks that have explicitly explored this domain. Asswad et al.[asswad2021data] highlighted the importance of protecting data ownership and reviewed representative related techniques. Chandrasekaran et al.[chandrasekaran2021sok] provided a comprehensive study on machine learning governance, including data ownership in machine learning. Recently, Du et al. [du2025sok] presented the first systematic SoK regarding dataset auditing, primarily focusing on surveying existing solutions. From the broader perspective of AI copyright protection, the majority of existing work focuses on watermarking techniques for models [ren2024copyright, shao2025sok], generated content [ren2024sok, zhao2024sok], and traditional multimedia [liu2024survey]. Lukas et al. [lukas2022sok] presented a toolbox to evaluate the robustness of DNN watermarking, which is fundamentally different from dataset auditing.
Compared with existing work, our paper makes the following four distinct contributions to the field. (1) This paper focuses on a specific scenario ($i.e.$, auditing the usage of a dataset) where existing techniques to watermark other objects are not feasible. The dataset is specifically used for training a DL model, and the model is the audit target. (2) This paper aims to provide a novel adversarial angle for evaluating existing dataset auditing methods. (3) Existing works fail to provide a comprehensive benchmark and evaluation regarding the SOTA dataset auditing methods. (4) This paper investigates new attack vectors, such as forgery attacks against dataset auditing, which are rarely discussed in prior papers [zhu2025evading].
6.2 Dataset Auditing for Generative Models
While dataset auditing has been actively explored for classification models, auditing generative models has only recently started to receive attention [qiu2025watermarking, hu2025membership, li2025towards] and remains largely underdeveloped. Many recent results suggest that even achieving stable results useful for auditing in real-world non-adversarial settings is extremely challenging [duan2023diffusion, duanmembership]. This difficulty arises from several inherent characteristics of generative models: (1) They are typically trained on massive-scale datasets (often containing billions of samples), where any individual data subset constitutes only a minuscule fraction, making both internal and external auditing signals significantly diluted. (2) The generation process is highly stochastic and sensitive to sampling seeds, further complicating reliable attribution. (3) Many real-world misuse cases manifest at the output level ($e.g.$, a generated image resembling a copyrighted artwork), where neither internal model parameters nor standard interfaces are available ($i.e.$, the box-free setting). In such cases, traditional auditing methods, which rely on model access or specific behavioral signatures, become ineffective or even inapplicable. While very recent works have achieved partial success in specific settings ($e.g.$, in fine-tuning scenarios [qiu2025watermarking, he2025towards, li2025towards]), designing stable and general-purpose auditing algorithms for generative models remains an open challenge in the community.
Given these complexities, existing auditing methods for generative models are not yet mature and may yield unstable or misleading results when evaluated under adversarial settings. Therefore, we caution that rushing into adversarial robustness evaluations in such an ill-defined setting without reliable baselines may misrepresent the true difficulty of the problem or misguide the community toward premature conclusions. Nevertheless, our framework and attack methodology—structured by three processing stages ($i.e.$, pre-processing, training, post-training) and three attack strategies ($i.e.$, decoupling, removal, detection)—are designed to be model-agnostic and generalize across different modalities. We leave the extension to generative models as an important direction for future work.
6.3 Potential Future Direction
Future Direction for Dataset Auditing. The demonstrated vulnerability of existing dataset auditing methods underscores an urgent need for fundamentally more robust and distinctive techniques. Future work may focus on developing novel auditing mechanisms resilient to these attacks. (1) For evasion, this could involve exploring IFs that are harder to decouple from the model utility. For EF-based methods, future works may seek EFs that are deeply integrated into the model’s core functionality rather than easily removable or detectable trigger patterns. A certified robust mechanism is also urgently needed. (2) For forgery, our empirical study suggests that complicating the auditing process may lead to better resistance. Consequently, exploring more complex methods, such as multi-bit watermarks [shao2025explanation] or cryptography [fiege1987zero], may be a promising future direction to improve dataset auditing.
Limitations and Future Directions for Evaluating Dataset Auditing. While DATABench provides a comprehensive benchmark, there are still some limitations. First, this paper primarily focuses on dataset auditing regarding image datasets since they are currently the most widely studied and feasible methods. There have been some preliminary works studying auditing datasets of other modalities, such as text [huang2024general, wei2024proving], point clouds [wei2024pointncbw], and audio [li2025cbw]. Providing a comprehensive evaluation of these works may be an important future work. Second, this paper does not cover all possible scenarios of dataset usage in DL. We primarily evaluate the two most important and widely adopted settings: training from scratch (Section 5) and fine-tuning (Appendix 5.4). Third, DATABench does not thoroughly explore finer-grained threat models, such as adversaries with varying levels of capabilities (an adversary with very limited computational resources or data). An exhaustive study of the attacker’s capabilities would be a meaningful direction for future work.
7 Conclusion
This paper presented a comprehensive adversarial evaluation of dataset auditing in DL systems, addressing the critical gap between existing auditing methods and their resilience against dedicated attacks. We introduced a novel internal feature and external feature taxonomy to classify auditing techniques, formally defined evasion and forgery attack objectives, and proposed systematic attack strategies to realize them. These contributions led to the development of DATABench, a benchmark featuring 17 evasion and 5 forgery attacks. DATABench also provides an extensible toolbox to design new dataset auditing methods and attacks. This toolbox can also facilitate the unified evaluation of auditing methods. Our empirical experiments of 9 representative auditing methods against DATABench revealed some critical findings, $e.g.$, none of them demonstrated sufficient robustness or distinctiveness under adversarial settings. These results highlight the significant vulnerability of current methods and underscore the urgent need for developing novel and adversarially resilient dataset auditing methods.
Appendix A Details of Dataset Auditing in DATABench
Shokri et al. (MIA) [shokri2017membership]. Shokri et al. proposed the first membership inference attack (MIA), leveraging models’ differential behavior ($e.g.$, higher confidence) on members versus non-members. The method trains multiple ‘shadow models’ to generate labeled prediction vectors (member/non-member), which then train a final classifier to distinguish target model members. We used 10 shadow models in our experiments.
Rapid [he2024difficulty]. Rapid is a SOTA MIA addressing inherent errors in difficulty calibration methods of reference-based attacks. It combines the original membership signal ($e.g.$, loss) with the calibrated score from reference models. Rapid trains a scoring model mapping both scores to a final membership prediction, learning to correct limitations of relying solely on calibration. We used 4 reference models in our experiments.
Dataset Inference (DI) [maini2020dataset]. DI assesses if a suspicious model shows higher prediction margins for the owner’s data points. It generates feature embeddings capturing this margin from the suspicious model using white-box (MinGD) or black-box (Blind Walk) methods. A confidence regressor, pre-trained on the owner’s model and data, scores these embeddings. Finally, a statistical hypothesis test aggregates scores to determine if the suspicious model used the owner’s dataset.
Data-use Auditing (DUA) [huang2024general]. DUA is a general dataset auditing framework. The owner creates two marked versions of each datum (preserving utility, maximizing difference), publishing one and hiding the other. Then, it uses a black-box Membership Inference (MI) score function and compares scores on the published versus hidden versions (contrastive MI). A sequential hypothesis test then determines if the published data was used for training.
DVBW [li2023black]. DVBW proposes the first and a general framework for backdoor-based dataset auditing. Specifically, DVBW utilizes poison-label backdoor attacks for dataset auditing and presents a hypothesis-test-based verification method. In our experiments, the target label is set to 1.
UBW-P/C [li2022untargeted]. UBW is proposed for harmless and stealthy dataset copyright protection using untargeted backdoors. UBW has two variants, UBW-P and UBW-C. UBW-P utilizes an untargeted poison-label backdoor attack that changes the labels of the poisoned samples randomly. In contrast, UBW-C adopts bi-level optimization to craft clean-label ($i.e.$, the labels of the watermarked samples do not change) samples. Finally, UBW also utilizes a hypothesis test to complete dataset auditing.
ZeroMark [guo2024zeromark]. ZeroMark is proposed for secure dataset auditing without disclosing watermarks. Inspired by the intrinsic property that boundary gradients of watermarked DNNs align with the watermark pattern, ZeroMark generates boundary versions of benign samples and calculates their boundary gradients under a label-only black-box setting. It then performs a hypothesis test based on the cosine similarity between the gradients and the secret watermark pattern to achieve auditing.
Domain Watermark (DW) [guo2023domain]. DW is proposed for harmless dataset auditing by identifying a hardly-generalized domain. Instead of causing misclassification, DW ensures watermarked models correctly classify specific ‘hard’ samples misclassified by benign models. DW employs bi-level optimization to generate the domain and craft visually indistinguishable clean-label modified samples for stealth.
Appendix B Details of Evasion Attacks in DATABench
Data Synthesis. Data synthesis involves generating artificial data, typically using generative models trained on the original dataset, to train the suspect model, thereby obscuring the direct link to the protected data. In our experiments, we use the DDIM model [ho2020denoising] as the generative model.
Denoising Filters. Denoising filters apply traditional image processing filters ($e.g.$, Gaussian, Median, Wavelet) to potentially remove noise-like external features embedded for auditing, effectively cleaning the data before model training. In our experiments, we set the kernel size to be 3.
Adversarial Training. Adversarial Training augments the training data with adversarially perturbed examples. We use two different implementations of adversarial training. AdvTraining-FGSM uses solely the FGSM method [goodfellow2014explaining] to generate adversarial examples while AdvTraining-Hybrid adopts three methods, FGSM [goodfellow2014explaining], PGD [madry2018towards], and C&W [carlini2017towards].
Adaptive Splitting Poisoned Dataset (ASD) [gao2023backdoor]. ASD is a training-time backdoor defense. ASD employs loss-guided splitting and meta-learning-inspired splitting to dynamically update two separate data pools. By leveraging the clean and poisoned data pools obtained through this adaptive partitioning, ASD effectively mitigates backdoors during training.
Noisy Output & Noisy Feature. Noise is injected into the model’s intermediate feature activations or final output predictions during inference to mask the subtle signals. In our experiments, we use Gaussian noise and the standard deviation is set to 0.5.
Reprogramming [chen2025refine]. The reprogramming technique modifies the model’s inference process, often using learnable input and output transformations, to neutralize the specific abnormal behaviors. In our experiments, we use a 10-layer U-Net as the input transformation module and train the U-Net for 10 epochs.
SCALE-UP [guo2023scale]. SCALE-UP is an effective black-box, input-level backdoor detection method. SCALE-UP identifies and filters malicious test samples by analyzing their prediction consistency during the pixel-wise amplification process.
Outlier Detection (OD). OD identifies input samples that are statistically different from typical data (potentially EF triggers), allowing the adversary to return a non-standard or random output to evade detection. In DATABench, we use two different models, $i.e.$, KNN and SVM.
To ensure simplicity and fairness in our evaluation, we adopt the most classical configurations for each attack. While this may introduce some degradation in model utility in certain scenarios, we note that such impacts can often be mitigated via parameter tuning, albeit potentially at the expense of attack efficacy. Users may further adjust these parameters based on their specific needs or evaluation preferences, which, however, falls beyond the scope of this work.
Appendix C Details of Forgery Attacks in DATABench
Appendix D Other Detailed Experimental Settings
As default settings, we train ResNet-18 on CIFAR-10 from scratch for in total 90 epochs with a learning rate of 0.03. The batch size is set to 128. Moreover, since batch normalization in ResNet-18 does not support DP-SGD, we alter these layers with modern group normalization. For EF-based methods, we use the Blended method [chen2017targeted], which applies a weighed blend of a specific trigger pattern and the original image, with the trigger pattern assigned a weight of 0.2.